# Social perception of faces in a vision-language model

### v1

@@ 1 Introduction
Humans spontaneously and quickly form opinions about someone’s trustworthiness, dominance, sincerity, intelligence, and other traits just by looking at their face in a photograph [2, 3, 4, 5]. Such “social perceptions” are not always accurate and yet can influence the outcome of social decision-making [6, 7, 8, 9, 10].
@@
We focus on CLIP [11], a state-of-the-art open-source VLM that is widely used for zero-shot classification [11], image retrieval [12], and guiding generative text-to-image models [13, 14, 15, 16]. CLIP models are trained on large vision-language-pair datasets [e.g., 17] gathered from the internet.
@@
We next ask whether CLIP’s social judgments are related to human ones and, if so, whether such judgments may mirror some of the common human biases. Forms of social bias in CLIP have been previously documented [12, 18, 19, 20, 21], as has bias in CLIP-guided generative text-to-image models (e.g., Stable Diffusion) [22, 23, 24, 25, 26]. For example, input-output bias has been measured by systematically prompting the model with language (such as “show me a CEO”) and then analyzing the demographics of the generated or retrieved image output.
@@
Our first innovation is using text prompts that are best at representing social perceptions based on evidence from social psychology. Previous studies use text prompts selected ad-hoc by the authors, e.g. denoting occupations or social roles [29, 30, 31]. In contrast, we rely on decades of social psychology literature and use words that have emerged as the consensus to represent the major ways in which people categorize each other. This literature investigates the semantic dimensions that perceivers use to predict the character and intentions of other people. Some attributes are of greater importance for effectively coordinating social behavior than others and thus serve as fundamental dimensions of social perception [32, 5]. According to the stereotype content model (SCM; [33]), the most relevant criteria are the persons’ intentions (Warmth) and their ability to carry out their plans (Competence). A more recent variant of the SCM model subdivided the two traits into three, named “Communion” (akin to Warmth), “Agency” (socio-economic success, akin to Competence), and a third dimension of political ideology “Beliefs” (progressive–conservative) [32]. We probe CLIP using text prompts expressing the content of both theories to test the robustness of results across them. These psychologically validated stereotype prompts have been used to study bias in AI language algorithms [34, 35, 36, 37, 38]. We introduce them to the study of bias in VLMs.
Figure 1: Overview of our method. We study social perception of human faces in CLIP, a vision-language model. Given one face image and a text prompt including a socially relevant attribute, we infer their CLIP embeddings and compute their cosine similarity within their shared latent space. Images are culled from three datasets: FairFace, UTKFace, CasualFace. FairFace and UTKFace are real-world (wild-collected) datasets which are annotated for legally protected variables, such as gender, race, and age. CausalFace is a synthetic dataset created by experimentally varying both protected (orange) and non-protected (purple) attributes. Text prompts are constructed based on attributes sourced from two theories quantifying social perception in humans (green box, bottom-right). To illustrate, a sub-trait of “warmth” is measured by the text prompt of “a photo of a friendly person” (friendliness is a warmth-associated word according to the Stereotype Content Model [39]).
@@ 2.1 Measuring social perception
We assess social perception in VLMs by computing the cosine similarities between text embeddings and image embeddings, following CLIP’s image retrieval metric [11]. Using this metric, we calculate the similarity between a set of image embeddings ${\bm{I}}$ (e.g., all face images of a demographic group) and a set of words $D$ that represent a social perception dimension (see Tab. S.2) as
@@
where ${\bm{t}}(d,t)$ is the text embedding of a prompt obtained by splicing a social trait adjective $d$ into a prompt template $t$ out of a predefined set of prompts $T$ (e.g., “a photo of a $<$adjective$>$ person”, see Tab. S.1 for all templates). Averaging similarities over a set of templates is a common practice to make results more robust [44].
@@
We introduced this metric because we observed a bias between demographic groups when using neutral text prompts. When we consider the distributions of cosine similarities of CausalFace images using a neutral text prompt (e.g., “a photo of a person”), we find that Asian women have significantly higher and Black men significantly lower cosine similarities compared to all other groups (p <=<=< = 0.01). Subtracting neutral prompt cosine similarities normalizes away this “ground” bias and more clearly reveals social biases (see Appx. G for more details).
@@ 2.2 Image datasets
Third, each prototype is varied in steps along age (+9 additional variations), smiling (+9), lighting (+7), and pose (+4). Consequently, CausalFace contains 30 (1 initial plus 29 variations) face images per prototype. Those variations are performed for each of the six gender-race demographic groups. Therefore, we obtain 180 face images per seed [^2] This number is larger than the 120 facial images per seed that Liang et al. [42] used for their analysis on bias in face recognition.. Overall, we used images from 100 different seeds, six prototypes per seed, and 30 variations per prototype. Consequently, our study uses 18,000 unique images. Additional details, including example images of the data set, can be found in Appx. E.2.
The important property of this dataset is that face attributes are manipulated systematically and independently, enabling an experimental approach and supporting causal interpretation of the results. To achieve this, the authors visually inspected whether manipulating one attribute induces unwanted changes in other attributes. Additionally, for each image, the authors collected nine human annotations per attribute to ensure that the apparent ID of the face was not changing and to check that the manipulations of the various attributes were successful. [42, Section 3.5 and Fig 4.]. The averages of the nine rater responses suggest that the manipulations are equally effective across demographic groups, i.e., no manipulation bias was found in CasualFace [42, Fig. 3]. The annotators were recruited worldwide through Amazon Mechanical Turk, and thus, the base of annotators was diverse.
Generating synthetic datasets that probe finer aspects of facial appearance, e.g. the differences between Mediterranean and Scandinavian physiognomies within the White group, is conceptually possible and may yield additional insights in both CLIP and human social judgments. In generating synthetic faces, one must, of course, pay attention to potential correlates between the variables of interest. For instance, pose and image color are known to confound gender bias [46]. The authors of CausalFace control for unwanted correlates through human annotation of their synthetic faces. They directly address pose and do not investigate image color. Our own examination reveals small differences in brightness between gender-race prototypes (Appx. F). Future research is needed to test for this potential confounder and, if necessary, correct any biases.
We used two popular real-world wild-sampled image datasets to compare our findings on CasualFace with findings from conventional observational studies.
@@
Similarly, UTKFace [48] images are annotated with demographic details such as race, gender, and age. From its 23,708 face images, we used a subset of 14,630 images applying the same criteria as for FairFace. Fig. S.7 shows example visualizations for both datasets.
@@ 3 Results
CausalFace images look very realistic. However, they are artificial, and it is reasonable to be concerned that they may differ from photographs of real people in ways that may affect our analysis. Therefore, we explored whether the statistical properties of CausalFace CLIP embeddings differ systematically from those of photographs of real people. To this end, we compared CausalFace to two observational datasets, FairFace [47] and UTKFace [48], with respect to six commonly used metrics: mean cosine similarities, markedness [49], WEAT [50], skew@k, max skew@k, and NDKL [21]. We find that the CausalFace measure of these six metrics is similar to those in datasets of real images (Tab. S.3). This gives us confidence that the CasualFace images are appropriate for our study. Detailed descriptions of these bias metrics can be found in Appx. B.
@@
Furthermore, smiling introduces more variation than age (Wilcoxon Rank-Sum test, independent samples, $p<0.001$), and lighting introduces as much variation as age (Wilcoxon Rank-Sum test, independent samples, $p=0.06$). In other words, protected and non-protected attributes cause comparable amounts of variation. Thus, considering a wide spectrum of protected and non-protected variables is necessary to understand and measure biases comprehensively. CausalFace is unique in providing the necessary attribute annotations.
@@
Previously, we explored the impact of non-protected variables at an aggregate level, averaging across all dimensions of social perception. Next, we shift the focus to a more detailed analysis. Varying lighting conditions and head orientation do not exhibit a clear trend: Higher $\Delta$ cosine similarities are observed under flat lighting, intermediate values for natural lighting, and lower values for unusual lighting. Similarly, for pose, higher $\Delta$ cosine similarities are found for extreme poses and lower values for central poses. For both variables, negative and positive valences show similar patterns and are positively correlated ($r_{\mathrm{lighting}}=0.93$, $r_{\mathrm{pose}}=0.79$, Fig. S.1 center and right panel ). This suggests that changing angles of light incidence (lighting) and head orientations (pose) do not significantly alter social perception but instead affect “legibility”: the ease with which different social perception dimensions can be interpreted or distinguished.
Conversely, varying facial expressions show a clear trend: As a face transitions from a frowning expression to a smile, it is perceived as more positive, more Progressive (Belief), and less negative. In other words, the opposing valences are negatively correlated ($r_{\mathrm{smiling}}$=-0.21, Fig. S.1 left panel). The finding that smiling positively influences social perception has two interpretations: First, it should be noted that facial expressions are unlabeled in widely used observational datasets like FairFace and UTKFace. Facial expression—and other non-protected variables—may correlate with protected attributes; therefore, ignoring such variables may compromise conclusions from observational studies on algorithmic bias. Second, it suggests that CLIP exhibits human-like social perception. CLIP can classify faces by race and gender (Fig. S.14), demonstrating its ability to make broad associations related to human perception. The just presented results go one step further and uncover CLIP’s ability to perform fine-grained social judgments.
@@
Figure 5: CausalFace shows different trait trajectories vis-a-vis age as compared to FairFace and UTKFace. Markers depict social perceptions of different intersectional groups. A second-degree polynomial is fitted for each group. Social perception is quantified as $\Delta$ cosine similarity, highlighting the difference in similarity of a valenced-text prompt from a neutral prompt. When testing images in CausalFace, Black women, highlighted by a red line, display distinct patterns: a U-shape trajectory instead of an inverse U-shape (Warmth) or extreme values (Warmth, Competence). Black women are perceived as most competent when young, which sharply and consistently decreases until old age. These trends are far less clear from FairFace (mid-row) and UTKFace (bottom-row). Additional social perception dimensions across age are plotted in Figs. S.3, S.4, S.5.
@@
Some of the specific age-related patterns we find diverge from human subject ratings. One study reported a decline in perceived Warmth for White women transitioning from young to middle-aged adulthood [53, p.8] – we find a similar decrease for Black women but not for White women. The same study found no change in Warmth perception for aging White men, whereas we observed an increase for older men across all three races [53, p.8].
@@ 4 Discussion
We find that CLIP exhibits human-like social perception. Similarly to human observers, it will classify faces according to race and gender, and increased smiling leads to more positive social perception. It is sensitive to age and relatively insensitive to viewpoint and lighting. CLIP thus mirrors humans in drawing trait inferences from face images [54]. Parallels between human and AI perception have been pointed out for Large Language Models (LLMs). They include similarities between word embeddings and human semantic memory [55], including group stereotypes [56, 57, 58, 59]. However, unlike LLMs which capture meaning from language, an already symbolic description, CLIP works with raw images where meaning is implicit in the pixel brightness values. While previous studies have shown that vision-language models (VLMs) can replicate certain human perceptions, such as susceptibility to visual illusions [60], our study is the first to directly demonstrate that these models can also reflect human-like social perceptions and stereotypes.
@@
Second, generative AI offers the opportunity to produce good quality synthetic stimuli and thus will help researchers move from observational to experimental approaches. An experimental approach offers considerable advantages for testing both AI systems and human subjects: controlled, systematic, incremental manipulation of variables across a large number of gender-race-age combinations allowed us to discover the new phenomena just described. Additionally, our conclusions are causal, thanks to mutually independent control of the variables of interest. For instance, varying facial expressions of Black women elicit the most extreme changes in social perceptions in CLIP as compared to other gender-race-intersections. This bias pattern had not been previously observed, likely because variations from wild-collected datasets are perturbed by noise from uncontrolled confound variables, and thus, some effects are hidden.
Our study is relevant to social psychology in we find that VLMs can generate testable hypotheses, enabling more efficient piloting and targeting of resources for human studies. A body of growing literature shows that LLMs replicate human behavior [66, 67] and thus are increasingly used to replace humans in behavioral studies [66, 68] and/or generate testable hypotheses [66, 67, 69]. Our study shows that this is also viable for VLMs.
@@ Acknowledgements
Appendix A Implementation details
All results were retrieved using CLIP ViT-B/32 as provided by OpenAI through the openai-clip Python library. The calculation of the cosine similarity of a set of images $I$ with a text dimension $D$ is described in Eq. 1. Tab. S.1 shows the full set of contextualized prompt templates used in this study. The dimensions sourced from theories of social perception used in this study are shown in Tab. S.2. We analyze the embeddings of the attributes listed in Tab. S.2 in Appx. D.
Table S.1: Prompt templates used for inferring text embeddings.
A photo of a $<$adjective$>$ person.
A $<$adjective$>$ person.
This is a $<$adjective$>$ person.
Cropped face photo of a $<$adjective$>$ person.
Table S.2: Adjectives per dimension. The Stereotype Content Model (SCM) and the ABC model both specify individual adjectives per dimension.
SCM ABC Warmth Competence Pos. Agency Neg. Agency Progressive Belief Conservative Belief Pos. Communion Neg. Communion warm competent powerful powerless science-oriented religious trustworthy untrustworthy trustworthy intelligent high-status low-status alternative conventional sincere dishonest friendly skilled dominating dominated liberal conservative friendly unfriendly honest efficient wealthy poor modern traditional benevolent threatening likeable assertive confident meek likable unpleasant sincere confident competitive passive altruistic egoistic
Appendix B Conventional bias metrics
Since CausalFace contains artificially generated images, one could question the dataset’s representativeness, as these faces might have different properties w.r.t. bias compared to real photos. Therefore, in the following, we compare CausalFace to the two observational datasets concerning six metrics: WEAT score, markedness, Skew@k, MaxSkew@k, and NDKL.
B.1 Adapted SC-WEAT
The traditional WEAT (word embedding association test) [70] requires opposite pairs of target items. Although these pairs have been validated in human linguistics, their antonymic relationship in CLIP’s embedding space remains unconfirmed. Given this uncertainty, we opted for the SC-WEAT, which obviates the need to define such pairs. In the following, we draw upon [50] who adapts the WEAT for investigating vision-language cosine similarities. More specifically, we adapt the Single-Category WEAT (SC-WEAT) as follows: Let ${\bm{A}}$ and ${\bm{B}}$ be two sets of image embeddings of equal size (e.g., men and women), and let ${\bm{d}}\in{\bm{D}}$ be a single text embedding from a certain social perception dimension ${\bm{D}}$ (e.g., Warmth). Then, we can calculate the test statistic for single text embedding as follows:

### v2

@@ 1 Introduction
Humans spontaneously and quickly form opinions about someone’s trustworthiness, dominance, sincerity, intelligence, and other traits just by looking at their face in a photograph [2, 3, 4, 5]. Such “social perceptions” are not always accurate and yet can influence the outcome of social decision-making [6, 7, 8, 9, 10, 11].
@@
We focus on CLIP [12], a state-of-the-art open-source VLM that is widely used for zero-shot classification [12], image retrieval [13], and guiding generative text-to-image models [14, 15, 16, 17]. CLIP models are trained on large vision-language-pair datasets [e.g., 18] gathered from the internet.
@@
We next ask whether CLIP’s social judgments are related to human ones and, if so, whether such judgments may mirror some of the common human biases. Forms of social bias in CLIP have been previously documented [13, 19, 20, 21, 22, 23], as has bias in CLIP-guided generative text-to-image models (e.g., Stable Diffusion) [24, 25, 26, 27, 28]. For example, input-output bias has been measured by systematically prompting the model with language (such as “show me a CEO”) and then analyzing the demographics of the generated or retrieved image output.
@@
Our first innovation is using text prompts that are best at representing social perceptions based on evidence from social psychology. Previous studies use text prompts selected ad-hoc by the authors, e.g. denoting occupations or social roles [31, 32, 33]. In contrast, we rely on decades of social psychology literature and use words that have emerged as the consensus to represent the major ways in which people categorize each other. This literature investigates the semantic dimensions that perceivers use to predict the character and intentions of other people. Some attributes are of greater importance for effectively coordinating social behavior than others and thus serve as fundamental dimensions of social perception [34, 5]. According to the stereotype content model (SCM; [35]), the most relevant criteria are the person’s intentions (Warmth) and their ability to carry out their plans (Competence). A more recent variant of the SCM model subdivided the two traits into three, named “Communion” (akin to Warmth), “Agency” (socio-economic success, akin to Competence), and a third dimension of political ideology “Beliefs” (progressive–conservative) [34]. We probe CLIP using text prompts expressing the content of both theories to test the robustness of results across them. These psychologically validated stereotype prompts have been used to study bias in AI language algorithms [36, 37, 38, 39, 40]. We introduce them to the study of bias in VLMs.
Figure 1: Overview of our method. We study social perception of human faces in CLIP, a vision-language model. Given one face image and a text prompt including a socially relevant attribute, we infer their CLIP embeddings and compute their cosine similarity within their shared latent space. Images are culled from three datasets: FairFace, UTKFace, CausalFace. FairFace and UTKFace are real-world (wild-collected) datasets which are annotated for legally protected variables, such as gender, race, and age. CausalFace is a synthetic dataset created by experimentally varying both protected (orange) and non-protected (purple) attributes. Text prompts are constructed based on attributes sourced from two theories quantifying social perception in humans (green box, bottom-right). To illustrate, a sub-trait of “warmth” is measured by the text prompt of “a photo of a friendly person” (friendliness is a warmth-associated word according to the Stereotype Content Model [41]).
@@ 2.1 Measuring social perception
We assess social perception in VLMs by computing the cosine similarities between text embeddings and image embeddings, following CLIP’s image retrieval metric [12]. Using this metric, we calculate the similarity between a set of image embeddings $\bm{I}$} (e.g., all face images of a demographic group) and a set of words $D$ that represent a social perception dimension (see Tab. 2) as
@@
where ${\bm{t}}(d,t)$ is the text embedding of a prompt obtained by splicing a social trait adjective $d$ into a prompt template $t$ out of a predefined set of prompts $T$ (e.g., “a photo of a <<adjective>> person”, see Tab. 1 for all templates). Averaging similarities over a set of templates is a common practice to make results more robust [46].
@@
We introduced this metric because we observed a bias between demographic groups when using neutral text prompts. When we consider the distributions of cosine similarities of CausalFace images using a neutral text prompt (e.g., “a photo of a person”), we find that Asian women have significantly higher and Black men significantly lower cosine similarities compared to all other groups (p <=<= 0.01). Subtracting neutral prompt cosine similarities normalizes away this “ground” bias and more clearly reveals social biases (see Appx. E for more details).
All results were retrieved using CLIP ViT-B/32 as provided by OpenAI through the openai-clip Python library.
Tab. 1 shows the full set of contextualized prompt templates used in this study. The dimensions sourced from theories of social perception used in this study are shown in Tab. 2.
Table 1: Prompt templates used to infer text embeddings.
A photo of a <<adjective>> person.
A <<adjective>> person.
This is a <<adjective>> person.
Cropped face photo of a <<adjective>> person.
Table 2: Adjectives per Dimension. The Stereotype Content Model (SCM) and the ABC model specify a set of psychometrically validated adjectives for each dimension.
SCM ABC
Warmth Competence Pos. Agency Neg. Agency Progressive Belief Conservative Belief Pos. Communion Neg. Communion
warm competent powerful powerless science-oriented religious trustworthy untrustworthy
trustworthy intelligent high-status low-status alternative conventional sincere dishonest
friendly skilled dominating dominated liberal conservative friendly unfriendly
honest efficient wealthy poor modern traditional benevolent threatening
likeable assertive confident meek likable unpleasant
sincere confident competitive passive altruistic egoistic
@@ 2.2 Image datasets
Third, each prototype is varied in steps along age (+9 additional variations), smiling (+9), lighting (+7), and pose (+4). Consequently, CausalFace contains 30 (1 initial plus 29 variations) face images per prototype. Those variations are performed for each of the six gender-race demographic groups. Therefore, we obtain 180 face images per seed [^2] This number is larger than the 120 facial images per seed that Liang et al. [44] used for their analysis on bias in face recognition.. Overall, we used images from 100 different seeds, six prototypes per seed, and 30 variations per prototype. Consequently, our study uses 18,000 unique images. Additional details, including example images of the data set, can be found in Appendix C
The important property of this dataset is that face attributes are manipulated systematically and independently, enabling an experimental approach and supporting causal interpretation of the results. To achieve this, the authors visually inspected whether manipulating one attribute induces unwanted changes in other attributes. Additionally, for each image, the authors collected nine human annotations per attribute to ensure that the apparent ID of the face was not changing and to check that the manipulations of the various attributes were successful. [44, Section 3.5 and Fig 4.]. The averages of the nine rater responses suggest that the manipulations are equally effective across demographic groups, i.e., no manipulation bias was found in CausalFace [44, Fig. 3]. The annotators were recruited worldwide through Amazon Mechanical Turk, and thus, the base of annotators was diverse.
Generating synthetic datasets that probe finer aspects of facial appearance, e.g. the differences between Mediterranean and Scandinavian physiognomies within the White group, is conceptually possible and may yield additional insights in both CLIP and human social judgments. In generating synthetic faces, one must, of course, pay attention to potential correlates between the variables of interest. For instance, pose and image color are known to confound gender bias [48]. The authors of CausalFace control for unwanted correlates through human annotation of their synthetic faces. They directly address pose and do not investigate image color. Our own examination reveals small differences in brightness between gender-race prototypes (Appendix D). Future research is needed to test for this potential confounder and, if necessary, correct any biases.
We used two popular real-world wild-sampled image datasets to compare our findings on CausalFace with findings from conventional observational studies.
@@
Similarly, UTKFace [50] images are annotated with demographic details such as race, gender, and age. From its 23,708 face images, we used a subset of 14,630 images applying the same criteria as for FairFace. Additional details on image selection for both datasets can be found in Appx. C.
@@ 3 Results
CausalFace images look very realistic. However, they are artificial, and it is reasonable to be concerned that they may differ from photographs of real people in ways that may affect our analysis. Therefore, we explored whether the statistical properties of CausalFace CLIP embeddings differ systematically from those of photographs of real people. To this end, we compared CausalFace to two observational datasets, FairFace [49] and UTKFace [50], with respect to six commonly used metrics: mean cosine similarities, markedness [51], WEAT [52], skew@k, max skew@k, and NDKL [22]. We find that the CausalFace measure of these six metrics is similar to those in datasets of real images (Tab. S.1). This gives us confidence that the CausalFace images are appropriate for our study. Detailed descriptions of these bias metrics can be found in Appendix A.
@@
Furthermore, smiling introduces more variation than age ($p<0.001$), and lighting introduces as much variation as age ($p=0.06$). In other words, protected and non-protected attributes cause comparable amounts of variation. Thus, considering a wide spectrum of protected and non-protected variables is necessary to understand and measure biases comprehensively. CausalFace is unique in providing the necessary attribute annotations.
@@
Previously, we explored the impact of non-protected variables at an aggregate level, averaging across all dimensions of social perception. Next, we shift the focus to a more detailed analysis. Varying lighting conditions and head orientation do not exhibit a clear trend: Higher $\Delta$ cosine similarities are observed under flat lighting, intermediate values for natural lighting, and lower values for unusual lighting. Similarly, for pose, higher $\Delta$ cosine similarities are found for extreme poses and lower values for central poses. For both variables, negative and positive valences show similar patterns and are positively correlated ($r_{\mathrm{lighting}}=0.93$, $r_{\mathrm{pose}}=0.79$, Fig. S.1 center and right panel).
This suggests that changing angles of light incidence (lighting) and head orientations (pose) do not significantly alter social perception but instead affect “legibility”: the ease with which different social perception dimensions can be interpreted or distinguished.
Conversely, varying facial expressions show a clear trend: As a face transitions from a frowning expression to a smile, it is perceived as more positive, more Progressive (Belief), and less negative. In other words, the opposing valences are negatively correlated ($r_{\mathrm{smiling}}$=-0.21, Fig. S.1 left panel).
The finding that smiling positively influences social perception has two interpretations: First, it should be noted that facial expressions are unlabeled in widely used observational datasets like FairFace and UTKFace. Facial expression—and other non-protected variables—may correlate with protected attributes; therefore, ignoring such variables may compromise conclusions from observational studies on algorithmic bias. Second, it suggests that CLIP exhibits human-like social perception. CLIP can classify faces by race and gender (Fig. S.12), demonstrating its ability to make broad associations related to human perception. The just presented results go one step further and uncover CLIP’s ability to perform fine-grained social judgments.
@@
Figure 5: CausalFace shows different trait trajectories vis-a-vis age as compared to FairFace and UTKFace. Markers depict social perceptions of different intersectional groups. A second-degree polynomial is fitted for each group. Social perception is quantified as $\Delta$ cosine similarity, highlighting the difference in similarity of a valenced-text prompt from a neutral prompt. When testing images in CausalFace, Black women, highlighted by a red line, display distinct patterns: a U-shape trajectory instead of an inverse U-shape (Warmth) or extreme values (Warmth, Competence). Black women are perceived as most competent when young, which sharply and consistently decreases until old age. These trends are far less clear from FairFace (mid-row) and UTKFace (bottom row). Additional social perception dimensions across age are plotted in Figs. S.3, S.4, and S.5.
@@
Some of the specific age-related patterns we find diverge from human subject ratings. One study reported a decline in perceived Warmth for White women transitioning from young to middle-aged adulthood [55, p.8]—we find a similar decrease for Black women but not for White women. The same study found no change in Warmth perception for aging White men, whereas we observed an increase for older men across all three races [55, p.8].
@@ 4 Discussion
We find that CLIP exhibits human-like social perception. Similarly to human observers, it will classify faces according to race and gender, and increased smiling leads to more positive social perception. It is sensitive to age and relatively insensitive to viewpoint and lighting. CLIP thus mirrors humans in drawing trait inferences from face images [56]. Parallels between human and AI perception have been pointed out for Large Language Models (LLMs). They include similarities between word embeddings and human semantic memory [57], including group stereotypes [58, 59, 60, 61]. However, unlike LLMs, which capture meaning from language, an already symbolic description, CLIP works with raw images where meaning is implicit in the pixel brightness values. While previous studies have shown that vision-language models (VLMs) can replicate certain human perceptions, such as susceptibility to visual illusions [62], our study is the first to directly demonstrate that these models can also reflect human-like social perceptions and stereotypes.
@@
Second, generative AI offers the opportunity to produce good-quality synthetic stimuli and thus will help researchers move from observational to experimental approaches. An experimental approach offers considerable advantages for testing both AI systems and human subjects: controlled, systematic, incremental manipulation of variables across a large number of gender-race-age combinations allowed us to discover the new phenomena just described. Additionally, our conclusions are causal, thanks to mutually independent control of the variables of interest. For instance, varying facial expressions of Black women elicit the most extreme changes in social perceptions in CLIP as compared to other gender-race intersections. This bias pattern had not been previously observed, likely because variations from wild-collected datasets are perturbed by noise from uncontrolled confound variables, and thus, some effects are hidden.
Our study is relevant to social psychology because it demonstrates that VLMs can generate testable hypotheses, thereby enabling more efficient piloting and targeted resource allocation for human studies. A body of growing literature shows that LLMs replicate human behavior [68, 69] and thus are increasingly used to replace humans in behavioral studies [68, 70] and/or generate testable hypotheses [68, 69, 71]. Our study shows that this is also viable for VLMs.
@@ Acknowledgements
Appendix A Conventional bias metrics
Since CausalFace contains artificially generated images, one could question the dataset’s representativeness, as these faces might have different properties with respect to bias compared to real photos. Therefore, in the following, we compare CausalFace to the two observational datasets concerning six metrics: WEAT score, markedness, Skew@k, MaxSkew@k, and NDKL.
A.1 Adapted SC-WEAT
The traditional WEAT (word embedding association test) [60] requires opposite pairs of target items. Although these pairs have been validated in human linguistics, their antonymic relationship in CLIP’s embedding space remains unconfirmed. Given this uncertainty, we opted for the SC-WEAT, which obviates the need to define such pairs. In the following, we draw upon Steed and Caliskan [52], who adapt the WEAT for investigating vision-language cosine similarities. More specifically, we adapt the Single-Category WEAT (SC-WEAT) as follows: Let $\bm{A}$} and $\bm{B}$} be two sets of image embeddings of equal size (e.g., men and women), and let ${\bm{d}}\in{\bm{D}}$ be a single text embedding from a certain social perception dimension $\bm{D}$} (e.g., Warmth). Then, we can calculate the test statistic for single text embedding as follows: